### v1

# Spiking Manifesto

Abstract
Practically everything computers do is better, faster, and more power-efficient than the brain. For example, a calculator crunches numbers more energy-efficiently than any human. Yet AI models are a thousand times less efficient than the brain. These models use artificial neural networks (ANNs) and require GPUs for the multiplication of huge matrices. In contrast, spiking neural networks (SNNs) of the brain have no matrix multiplication and much smaller energy requirements. This manifesto proposes a framework for thinking about popular AI models in terms of spiking networks and polychronization, and for interpreting spiking activity as nature’s way of implementing look-up tables. This offers a way to convert AI models into a novel type of architecture with the promise of a thousandfold improvement in efficiency. Code is available at https://github.com/izhikevich/SNN
I The manifesto
Llama 3 has artificial neurons, one hundred billion (synaptic) parameters, and consumes tens of watts in inference mode. In contrast, the human brain has spiking neurons, one quadrillion synapses, and consumes only 20 watts. Scaled down 10,000-fold to the size of Llama 3, the brain would require only 2 mW – enough to power it for a month with a single AAA battery! There are two reasons for the power efficiency of the brain:
Efficient processing: Spiking neural networks (SNNs) of the brain have sparse activity. If the input to a spiking neuron is below its threshold, the neuron is silent and no signal is transmitted. Only when the input exceeds the threshold the neuron generates a brief all-or-none event, called spike in Fig. 1a, that propagates to target neurons. This signaling sparsity provides unparalleled efficiency. In contrast, artificial neurons in popular AI architectures generate continuous-valued neuronal activations. At each time step, activations are transmitted to target neurons and multiplied by synaptic weights to calculate new inputs, resulting in a massive demand for GPU matrix multiplications.
Figure 1: (a) The all-or-none spiking response of a cortical neuron to a random input. (b) 60 neurons, each firing only one spike, have a representation capacity 60!, which is greater than the number of particles in the visible universe.
Efficient encoding: Neuronal activations in ANNs are often interpreted as probabilities of spiking or as instantaneous firing rates. Both interpretations ignore the timing of spikes and miss the opportunity to have a combinatorially explosive representation capacity. In Fig. 1b, we depict 60 neurons, each firing only 1 spike during a certain time period. From an ANN’s point of view, this corresponds to all activations being equal to $1$, that is, to a single feature vector, a single embedding, with activation values $(1,1,\ldots,1)\in\mathbb{R}^{60}$. Changing the order of spikes does not change the activation values. However, from an SNN’s point of view, this is just one pattern out of $60$ factorial possible patterns of different relative spiking orders. 60! is greater than $10^{80}$ – the number of particles in the visible universe. Therefore, a single representation value from an ANN’s point of view has a practically infinite representation capacity from an SNN’s point of view.
Figure 2: Classification of ANNs and SNNs based on the ”memory vs. processing” (space vs. time) demand. SNNs typically have large memory footprint, but only a tiny fraction is used at any given forward pass, resulting in low memory bandwidth requirement.
High representation capacity requires less memory, and sparse activity requires less processing, putting SNNs in the opposite quadrant from ANNs in the memory vs. processing diagram in Fig. 2. While most SNN research emphasizes sparse activity, this manifesto calls to take advantage of the combinatorially explosive representation capacity necessary to achieve the top-right quadrant.
One could argue that the interpretations of neuronal activations do not matter because the states of both ANNs and SNNs can be described by vectors encoding features; indeed, the spike-timings in Fig 1b can be described by a 60-dimensional vector of spike latencies. However, the meaning we assign to these vectors fundamentally changes how we operate on them, summarized in Fig. 3. In ANNs, the interpretation of vector values as continuous activations utilizes relationships like orthogonality and correlation, and relies heavily on operations such as dot products and matrix multiplications. This computational paradigm is what allowed ANNs to win the GPU hardware lottery [2], as GPUs are highly optimized for these exact operations.
Figure 3: The distinction between ANNs and the SNN described in this manifesto. The linear capacity $e^{\varepsilon^{2}n}$ to represent $\varepsilon$-orthogonal vectors, given in Exercise 5.1.15 in [3], is dwarfed by the combinatorial capacity $n!$.
Figure 4: (a) Precise and repeatable sequence of spikes (with error bars) in in vivo recordings of 64 human anterior temporal lobe neurons while presenting visual images of places. (b) The same neurons polychronize with different repeatable sequences while the subject is viewing other image categories [8]. Syn-chronous – one timing; poly-chronous – multiple timings [9].
This manifesto treats spiking patterns as vectors of latencies. This interpretation makes orthogonality, correlations, and matrix multiplications irrelevant. Instead, features are represented by polychronous spiking patterns, illustrated in Fig. 4, where the precise order of spikes within each pattern is important. Polychronization was first observed in biologically-detailed brain models [4, 5] and then in the human brain [6, 7, 8]. These patterns represent memories and their number is often greater than the number of synapses in the network [9]. However, simulating spiking networks requires modeling membrane potentials of individual spiking neurons and direct interactions between such neurons - a monumental task best suited for traditional neuromorphic hardware like TrueNorth [10], LOIHI [11], or SpiNNaker [12].
Figure 5: Spiking dynamics is treated as a transition from one spiking pattern $x^{l}$ to another spiking pattern $x^{l+1}$, where $l$ could denote different layers in a feedforward network or different cycles within the same layer. There is no time variable within each spiking pattern. Instead, it is described by the vector $x=(x_{1},\ldots,x_{n})\in\mathbb{R}^{n}$, where each $x_{i}$ is the latency (timing) of spike of the $i$th neuron relative to some time stamp 0.
The first pillar of this manifesto is to abstract away from individual neurons. We model the network transition as a direct map from latency vector to latency vector, i.e., from spiking patterns in one layer to spiking patterns in the next layer (or the same layer at the next periodic cycle), illustrated in Fig. 5. This is accomplished by detecting spiking patterns via look-up tables and using the tables to determine latencies in the next layer. Because of the continuity of latency vectors, we can differentiate them and use gradient descent for training. As a result, we can match the high performance of leading AI models, but implement them in a look-up table format that is inherently more efficient.
The second pillar of this manifesto is the claim that spiking networks are nature’s way of implementing look-up tables. Unlike matrix multiplication but similar to look-up tables, spiking networks store a vast number of parameters (synapses) yet use only a tiny fraction at any given time because most neurons are silent. Similarly to look-up tables, spiking networks determine the identity (index) of each fired neuron; and exactly like look-up tables, spiking networks utilize only the synapses from that fired neuron while all the other synapses remain untouched until needed.
The AI field has come a long way from the original view of neurons as gates in logical circuits (McCulloch-Pitts [13]) to the modern linear algebra view: neuronal states as vectors, transition from one state to another via matrix multiplication, and learning as adjustment of the weight matrices. It is time to take the analogous step and treat spiking patterns as the fundamental states of the network, the transition from one pattern to another via look-up tables, and learning as the adjustment of these look-up tables. This approach allows for the replacement of all matrix multiplication in deep networks and transformers with look-up tables, and it suggests a novel hardware architecture to accelerate these models, one that is distinct from GPUs and traditional neuromorphic systems [14].
We discuss relative capacity of ANNs and SNNs (Sect. VIII-B), the relationship to locality-sensitive hashing (Sect. VI), to mixture-of-experts (Sect. VIII-L), finite-state machines (Sect. VIII-M), decision trees, random forests, and ferns (Sect. VIII-N), spiking implementations of transformers via quantizations (Sect. VIII-O), and neuromorphic systems (Sect. VIII-P) in Appendix.
II The model
To exploit the massive combinatorial capacity of spiking patterns – the feature that gives our network its efficiency – we need a way to distinguish different patterns, even those that differ by the timing of a single spike. For this, we consider a special case when all neurons fire exactly one spike during a certain interval, as in Fig. 5. Any such spiking pattern can be described by a vector of latencies $x\in\mathbb{R}^{n}$, where $n$ is the number of neurons. That is, the $i$th neuron fires a spike at a time $x_{i}$ relative to the center of the interval, so that $x_{i}$ could be positive or negative. In the case of feedforward networks, we want to describe the transition from spike latencies in the layer $x$ to the spike latencies in the next layer $y$. In the case of recurrent networks, we describe the transition from $x$ at one temporal cycle to the same $x$ at the next cycle.
The transition from $x$ to $y$ in ANNs is done via multiplying $x$ by a matrix of synaptic weights and then applying a nonlinear function, such as ReLU. However, matrix multiplication distinguishes poorly between relative timings of individual spikes. A better way to detect spiking patterns is to compare pairwise latencies of neurons in $x$. This allows us to enumerate various patterns using discrete integer indices $j$. Then, we use these indices to select pre-stored synaptic vectors from look-up tables, which are then summed to form the output vector $y$, effectively translating one spiking pattern $x$ into another spiking pattern $y$ for the subsequent layer.
Figure 6: Various spiking patterns in layers 1 and 2 are represented by the vector of latencies (timings) $x,y\in\mathbb{R}^{10}$. Transition from $x$ to $y$ is done through $n_{t}=3$ look-up tables. Each look-up table monitors $n_{c}=3$ pairs of ”anchor” neurons, selected randomly during initialization. The first (red) look-up table monitors $(x_{1},x_{2})$, $(x_{2},x_{3})$, and $(x_{3},x_{5})$. At each step, it evaluates 3 conditions, if $x_{1}>x_{2}$, if $x_{2}>x_{3}$, and if $x_{3}>x_{5}$, each resulting in binary 0 or 11. Concatenating these 3 binary values results in an integer between 0 and $7$, e.g., $011_{\rm binary}=3$ in this figure (indexing starts with $000$). The integer is used as an index into the red table to retrieve the vector of synaptic values. These values are used as inputs to the vector of latencies $y$ in layer 2, resulting in various spiking patterns there.
Let us formalize this. The transformation from $x$ to $y$, illustrated in Fig. 6, involves $n_{t}$ look-up tables, each tracking $n_{c}$ ”anchor” neurons using the following steps:
  • Assign: During initialization, assign $n_{c}$ random pairs of input neurons (anchors) to each table.
  • Compare: During inference, each table makes pairwise comparisons of latencies $x$ of its $n_{c}$ anchor neurons.
  • Concatenate: The resulting $n_{c}$ bits are concatenated to form a unique binary index $j$.
  • Select: The index $j$ selects the row in the look-up table containing the vector of synaptic values to be used as an output from $x$ to the next layer $y$.
For example, Fig. 6 has $n_{t}=3$ tables, each performing $n_{c}=3$ comparisons, and hence having $2^{n_{c}}=8$ rows. Red table in the figure evaluates 3 conditions: $x_{1}>x_{2}$, $x_{2}>x_{3}$, and $x_{3}>x_{5}$, resulting in binary $011$ corresponding to the synaptic vector stored in row $j=3$ (row numbering starts with binary $000$). The entire row is used as the input for the next layer of neurons, $y$, achieving the desired transformation without any matrix multiplication. One can think of each look-up table as a sparse network of $2^{n_{c}}$ spiking ”detector” neurons, each tuned to its own order of input spikes of the anchor neurons. Every spiking pattern excites a single detector neuron.
Let vectors $a_{i}$ and $b_{i}$ denote the $n_{c}$ pairs of indices of the anchor neurons associated with the $i$th table. For example, the red table in the figure has $a_{1}=(1,2,3)$ and $b_{1}=(2,3,5)$. The $i$th table looks at the pairs $u_{ir}=x_{a_{ir}}-x_{b_{ir}}$ and evaluates the condition $u_{ir}>0$ to get $n_{c}$ binary values for all $r=1,\ldots,n_{c}$. Then, the row index $j$ is just concatenation of these binary values
$j=H_{i}(x)=\mbox{concat}(\ u_{i1}\!>\!0,\ \ldots,\ u_{in_{c}}\!>\!0\ ),$ (1)
which is a function of $x$. Each such index corresponds to a spiking pattern exhibited by anchor neurons with indices $a_{i}$ and $b_{i}$, and the table can distinguish $2^{n_{c}}$ such patterns.
Let $s_{ijk}$ denote the synaptic value in table $i$, row $j$ (i.e., corresponding to polychronous pattern $j$), and column $k$ (i.e., projecting to $y_{k}$). Since selecting index $j$ is a function of $x$, we can write $s_{iH_{i}(x)k}$, but this notation is cumbersome. Instead, for simplicity, we will write $s_{ixk}$, keeping in mind that $x$ stands for an appropriate second index $j$ of $s_{ijk}$, different for different $i$. Now, the transition from $x$ to $y$ can be written as
$y_{k}=\sum_{i=1}^{n_{t}}s_{ijk}\ \ \mbox{ where }\ \ j=H_{i}(x)$ (2)
or, inserting $j=H_{i}(x)$ into the index, as
$y_{k}=\sum_{i=1}^{n_{t}}s_{iH_{i}(x)k}\ \ \ \ \ \mbox{or }\ \ \ \ y_{k}=\sum_{i=1}^{n_{t}}s_{ixk}$
or using vector notation $S_{ix}=(s_{ix1},\ldots,s_{ixn})\in\mathbb{R}^{n}$ as
$y=\sum_{i=1}^{n_{t}}S_{ix}\ \ \ \ \mbox{or }\ \ \ \ y=S_{x}\;$ (3)
All these notations are equivalent. $S_{x}$ is just a shorthand to denote the sum of constant vectors. Variable $x$ only plays a role in determining which rows from each look-up table $S_{i}$ to use; see the first half of the pseudocode in Fig. 10.
Motivated by the dynamics of cortical neurons (see Sect. VIII-E), we use simple addition for the update from $x$ to $y$. Thus, it is a misnomer to call $s_{ijk}$ weights, since the term ”weights” implies that they are multiplied by something, whereas in spiking networks the synapses are always added to something. The Eq. (3) is nevertheless a nonlinear operation because of the selection of the index $j=H_{i}(x)$, i.e., selecting which $s_{ijk}$ to add. The combined action of summing the look-up tables is defined as the look-up transformation $S_{x}$ (LUT $S_{x}$). In the Appendix, we will discuss other possible choices of $H_{i}(x)$.
We now make two observations to justify Fig. 2:
  • Low processing demand. The conversion from a spiking pattern to a look-up table is just a bit concatenation. A look-up table with $n_{c}$ inputs distinguishes $2^{n_{c}}$ different patterns at no additional processing cost. A LUT with $n_{t}$ look-up tables has the capacity to distinguish $2^{n_{t}n_{c}}$ different spiking patterns.
  • Low memory demand. While memory footprint of each look-up table might be large, only one row is retrieved during each forward pass, resulting in low memory bandwidth demand. Memory storage is cheap; memory access is expensive.
The representation of a practically infinite set of all possible spiking patterns as a set of sparse indices of look-up tables is a form of locality-sensitive hashing [15], and it is arguably the most efficient method of encoding and retrieving information. We explore this relationship in Sect. VI.
III Gradients and error backpropagation
Learning through gradient descent involves partial derivatives of (3), which are discontinuous with respect to $x$. Indeed, $S_{ix}$ is piecewise constant unless $x_{a_{ir}}=x_{b_{ir}}$ for some $r$, in which case a small perturbation flips the order of spikes, changes the index $j=H_{i}(x)$, and flips $S_{ij}$ to some other value. This complication is related to the general problem of applying error backpropagation to all SNNs – small perturbations tweak timings of spikes, but if a neuron is near its threshold, perturbations can add or remove a spike, making the whole dynamics discontinuous. However, due to the special way we represent spiking patterns, there is a path to do error backpropagation through a look-up table.
Figure 7: Error backpropagation through look-up tables: Selection of the minimal anchor pairs, $u_{i}$, for the learning rule (5). Flipping the order of spikes in any such pair will result in a change in the index of the corresponding table. Notice that the smallest latency difference is between $x_{7}$ and $x_{8}$, but none of the look-up tables compares these two neurons and none will be affected by flipping their order.
III-A Surrogate gradient
For each look-up table, let us denote the anchor pair having the smallest absolute value as $u_{i}=x_{a_{i}}-x_{b_{i}}$, as in Fig. 7. Among all other latency differences monitored by table $i$, this $u_{i}$ requires the smallest perturbation of $x$ to flip its sign. This flipping corresponds to the index in the $i$th look-up table, $j=H_{i}(x)$, flipping one bit to become $\bar{j}=H_{i}(\bar{x})$, causing the synaptic vector $S_{ij}$ to flip to $S_{i\bar{j}}$, i.e., $S_{ix}$ to $S_{i\bar{x}}$. When $u_{i}\approx 0$, we are not sure whether we should use synaptic values $S_{ix}$ or flipped values $S_{i\bar{x}}$. And when $u_{i}=0$, probably the average of the two is the best choice. Let continuous symmetrical function $U(u)$ in Fig. 8 represent our ”uncertainty” about the choice. It should satisfy $U(u)\approx 0$ for large $|u|$ and $U(0)=0.5$ reflecting the biggest uncertainty.
Figure 8: Examples of symmetrical U-shaped ”uncertainty” functions to smoothen the transitions from one polychronous pattern to another.
Then, we can use $U$ to make the dynamics smooth, at least in the small neighborhood of $x$ that only includes flipping of signs of $u_{i}$. This is similar to the standard method of ”surrogate gradients” in spiking networks [17, 18]. We replace (3) with a surrogate function
$y=\sum_{i=1}^{n_{t}}\left\{S_{ix}+U(x_{a_{i}}-x_{b_{i}})(S_{i\bar{x}}-S_{ix})\right\}\;.$ (4)
Basically, as $u_{i}\rightarrow 0$, the synaptic values approach the average $(S_{ix}+S_{i\bar{x}})/2$. Other choices, e.g. applying $U$ to all pairs or ignoring the flipping, are discussed in Sect. VIII-K.
Differentiating this equation with respect to $x_{m}$ gives zero Jacobian vector $\partial y/\partial x_{m}=0$ unless $m$ equals either $a_{i}$} or $b_{i}$}, in which case
$\frac{\partial y\ }{\partial x_{a_{i}}}=-\frac{\partial y\ }{\partial x_{b_{i}}}=U^{\prime}(x_{a_{i}}-x_{b_{i}})(S_{i\bar{x}}-S_{ix})\;.$ (5)
Oppositely signed derivatives represent the two opposing forces that either pull $x_{a_{i}}$ and $x_{b_{i}}$ together or push them apart in Fig. 7. When neuron $x_{m}$ anchors the minimal comparison across multiple look-up tables, the derivatives are summed.
We only use $U$ in the training mode to calculate gradients, but we use (3) without $U$ in the inference mode. The choice of $U$ changes the training results very little. In simulations, we use $U(u)=0.5/(1+|u|)$ in Fig. 8, left. Notice that neither $S$ nor $x$ have any scale in the model. Using the nonlinear function $U$ introduces the scale, limiting the plasticity window to sufficiently small latency differences $u\approx 0$.
III-B Error backpropagation
Let us consider a generic multi-layer SNN of the form
$x^{l+1}=S^{l}_{x^{l}}\;,$ (6)
where $l$ is the layer number. Let $\mathcal{L}$ denote the loss (error) function that we want to minimize. Differentiating $\mathcal{L}$ with respect to each $x^{l}_{m}$ and applying the chain rule results in the standard recursive expression
$\frac{\partial\mathcal{L}\ }{\partial x^{l}_{m}}=\frac{\partial\mathcal{L}\ \ }{\partial x^{l+1}}\cdot\frac{\ \partial x^{l+1}\ }{\partial x^{l}_{m}}$
that shows how the error at layer $l+1$ propagates back to layer $l$. Here, the Jacobian vectors $\partial x^{l+1}/\partial x^{l}_{m}$ are either 0 or are given by the equation (5) with $y=x^{l+1}$. Introducing variable
$g^{l}_{i}=\frac{\partial\mathcal{L}\ \ }{\partial x^{l+1}}\cdot\left(S^{l}_{i\bar{x}^{l}}-S^{l}_{i{x}^{l}}\right)\;,$ (7)
we can write the gradient backpropagation equation as
$\frac{\partial\mathcal{L}\ }{\partial x^{l}_{a^{l}_{i}}}=-\frac{\partial\mathcal{L}\ }{\partial x^{l}_{b^{l}_{i}}}=U^{\prime}(x^{l}_{a^{l}_{i}}-x^{l}_{b^{l}_{i}})\;g^{l}_{i}\;.$ (8)
This learning rule, used in all simulations in this paper, has a simple interpretation: $g_{i}$ measures which synaptic vector, $S_{i\bar{x}}$ or $S_{i{x}}$, is better aligned with the gradient vector $\partial\mathcal{L}/\partial x^{l+1}$. If $g_{i}$ is negative, then keeping the current $S_{ix}$ is better and the minimal latency difference should be further away from zero. If $g_{i}$ is positive, then flipping is better and the minimal difference should be closer to zero. This has to be done for all look-up tables in all layers. A geometrical illustration of this rule is in Fig. 14.
Each table contributes to the gradient two equal numbers of the opposite sign, $\partial\mathcal{L}/\partial x^{l}_{a}$ and $\partial\mathcal{L}/\partial x^{l}_{b}$. As a result, the gradient vector $\partial\mathcal{L}/\partial x^{l}$ has zero mean, which causes the synaptic vectors, $S^{l}_{ix}$, and hence the latency vectors, $x^{l}$, to maintain zero mean, providing self-regularization and stability.
In Appendix, Sect. VIII-K, we explore variants of this rule. One variant takes into account all latency differences for all $n_{c}$ comparisons within each table. Another one considers only one table (with minimal $u_{i}$) in the entire layer to compute the gradient. That latter rule can be implemented without matrix multiplication, hinting on the possibility that the same spiking hardware can do both, training and inference.
IV Examples
To illustrate the versatility of our approach, we demonstrate the conversion of deep ANNs, recurrent networks (RNNs), and Transformers into the SNN form. We adhere to a single rule: preserve the model’s overall architecture, treat each activation vector $x$ as a latency vector, and simply replace all matrix multiplications with look-up table transformations. This process is akin to compiling high-level code to specialized assembly language, where the fundamental complexity is managed beneath the surface.
Figure 9: Replacing matrix multiplication with look-up transformation (3) converts a deep ANN into a deep SNN.
IV-A Deep SNN
Consider the simplest ANN – multi-layer perceptron (MLP) with residual (skip [19]) connections
$x^{l+1}=x^{l}+\mbox{ReLU}(M^{l}x^{l}+b^{l})\;.$ (9)
Here synaptic weight matrix $M^{l}\in\mathbb{R}^{n\times n}$ and bias vector $b^{l}\in\mathbb{R}^{n}$ are layer dependent, $l=0,\ldots,L$, and nonlinearity is conveyed via ReLU(u)=(u)= max(0,u)(0,u). This is a fundamental block in many ANN models, including the feedforward layer (FFN) in transformer architecture.
Figure 10: Pseudocode of the deep SNN (10).
1:Initialize look-up tables: zero or random synaptic vectors $S^{l}_{ij}$
2:Initialize anchors: random pairs $a^{l}_{ir}\neq b^{l}_{ir}$ for each table
3:Forward pass (inference):
[…]
5:for $l=0\textbf{ to }L$ do     // all layers
6:  $x^{l+1}\leftarrow x^{l}$   // residual connections
7:  for $i=1\textbf{ to }n_{t}$ do // all look-up tables
8:    $j\leftarrow 0$   // look-up table index
9:    for $r=1\textbf{ to }n_{c}$ do // all comparisons
10:     $u_{ir}\leftarrow x_{a_{ir}}-x_{b_{ir}}$ // anchor neurons
11:     $j\leftarrow\textbf{concat}(j,u_{ir}>0)$ // Eq. 1: determine the index
12:    end for
13:    cache $j$ and $u_{i},r_{i}$ corresponding to $\min(|u_{ir}|)$
14:    $x^{l+1}\leftarrow x^{l+1}+S^{l}_{ij}$   // Eq. 6: synaptic transmission
15:  end for
16:end for
17:output $\leftarrow x^{L+1}$
18:Backward pass (learning):
[…]
21:for $l=L\textbf{ down to }0$ do     // all layers backwards
22:  $v^{l}\leftarrow v^{l+1}$   // residual connections
23:  for $i=1\textbf{ to }n_{t}$ do   // all look-up tables
24:    $\textbf{uncache }j$ and $u_{i},r_{i}$   // corresp. to $\min(|u_{ir}|)$
25:    $\bar{j}\leftarrow j\text{ \bf xor }2^{r_{i}}$   // flip $r$th bit of $j$
26:    $g_{i}\leftarrow v^{l+1}\cdot(S^{l}_{i\bar{j}}-S^{l}_{ij})\,$ // Eq. 7: gradient alignment
27:    $v^{l}_{a}\leftarrow v^{l}_{a}+U^{\prime}(u_{i})\ g_{i}$ // Eq. 8 with $a=a^{l}_{ir_{i}}$
28:    $v^{l}_{b}\leftarrow v^{l}_{b}\,-U^{\prime}(u_{i})\ g_{i}$ // Eq. 8 with $b=b^{l}_{ir_{i}}$
29:    $S^{l}_{ij}\leftarrow S^{l}_{ij}-\varepsilon v^{l+1}$   // synaptic update (learning)
30:  end for
31:end for
Figure 9 illustrates our general principle: The overall architectural flow of vectors $x^{l}\in\mathbb{R}^{n}$ is preserved; we simply replace the standard ReLU and MatMul operations with look-up table transformations:
$x^{l+1}=x^{l}+S^{l}_{x^{l}}\;.$ (10)
Notice in the pseudocode in Fig. 10 that we do not need to keep activation vectors $x^{l}$ for backward (learning) pass. Instead, we rely on index caching (line 13). This index caching becomes crucial for the V-index caching in the efficient SNN implementation of attention in Sect. IV-C – the SNN’s equivalent of KV-caching in ANN transformers.
IV-B Spiking RNN
Recurrent neural networks (RNNs) serve as the foundation of many impactful early LLMs such as Long Short-Term Memory (LSTM) [20, 21] or Gated Recurrent Unit (GRU) [22]. At any given time step $t$, the state of the network, $h_{t}$, is determined by the input $z_{t}$ and the previous, ”hidden”, state $h_{t-1}$ via some transformation $h_{t}=f(h_{t-1},z_{t})$ that typically involves matrix multiplications. The simplest implementation of an RNN is an Elman network [23]:
$h_{t}=S_{h_{t-1}}+z_{t}\;,\ \ \ h_{0}=0\;,$ (11)
where LUT $S_{h}$ has the form (3). In addition, it also has a learnable look-up table $E$ that embeds the $t$th input character into $z_{t}\in\mathbb{R}^{n}$ and a learnable LUT $U_{h_{t}}$ of the form (3) that unembeds the hidden state $h_{t}$ into the output probability distribution over the vocabulary set. In Sect. VIII-F we discuss other variants of spiking RNNs.
IV-C SNN Transformer
Figure 11: (a) Replacing an ANN transformer with SNN transformer. Each (optional) Feed Forward network is a single LUT (12) and each attention head is the LUT (15) depicted in Fig. 12.
Now, we are ready to implement an SNN transformer. The first step, common to both ANNs and SNNs and depicted in Fig. 11, is to use an embedder look-up table to convert each input token into an embedding vector $z_{i}\in\mathbb{R}^{n}$, where $i$ denotes the position of the token in the input string of length $n_{inp}$. The reverse is done at the last step, projecting an embedding vector back to vocabulary space and then taking the softmax to get probability distribution of the output tokens. ANN transformer un-embeds via matrix multiplication while SNN transformer uses LUT of the form (3). Both, embedder and un-embedder are learnable. The middle of the transformer is the ”attention + feedforward” block repeated $N$ times. We denote inputs to the attention module as $z_{i}$ and outputs as $x_{i}$; the feedforward network (FFN) converts $x_{i}$ back to $z_{i}$.
The SNN implementation of FFN layer,
$z_{i}=x_{i}+S_{x_{i}}\,$, (12)
is just a single-layer LUT (10) with pseudocode in Fig. 10. We show later that this module is optional, owing to the inherent capabilities of SNN attention module.
The standard dot-product attention module,
$\mbox{softmax}(\frac{\ QK^{\top}}{\sqrt{d_{k}}})\,V\;,$ (13)
has a built-in multiplication of matrices, so obtaining $Q$, $K$, and $V$ via LUTs would still leave softmax bottleneck and matrix multiplication of $V$ by the softmax attention scores.
Figure 12: (a) Replacing a single head of transformer’s attention (13) with SNN (15). (b) Zoom-in on the concatenation and look-up transformation (LUT). Optional mask is omitted for simplicity.
Among many approaches discussed in Appendix, Sect. VIII-G, the most efficient one is presented in Fig. 12. It captures the long-distance dependencies between different input embeddings by concatenating input vectors
$z_{i},z_{j},PE_{i-j}]\in\mathbb{R}^{2n+p}$ (14)
for every Query-Key pair $z_{i}$ and $z_{j}$. Here, the positional encoder $PE_{i-j}\in\mathbb{R}^{p}$ is a learnable vector that encodes the relative distance from $z_{i}$ to $z_{j}$. If $p!>n_{inp}$, then $PE$ provides enough context to encode all relative positions as a pattern of spike latencies.
Each such concatenated vector is then used as an input to
$x_{i}=z_{i}+\sum_{j=1}^{i-1}V_{[z_{i},z_{j},PE_{i-j}]}\;,$ (15)
where LUT $V$ is the same for all embedding pairs, but different for different heads and layers. In self-attention, $z_{i}$ can only look at the earlier tokens, so we only consider $j<i$. Interestingly, including the ”self” term $j=i$ into the sum reduced the performance of the model. Outputs of multiple attention heads are added without any weighting, though one can treat each look-up table as its own attention head.
Notice that there is no softmax operation here. The attention network learns the relative magnitudes (attention scores) of value vectors $V_{[z_{i},z_{j},PE_{i-j}]}$ by being exposed to various pairs of embeddings at various locations. This is conceptually similar to how different attention heads learn different long-distance dependencies even though there is no direct competition between them. In Appendix, we also implemented the ”QK transformation” via LUTs to get attention scores $a_{ij}=A_{[z_{i},z_{j},PE_{i-j}]}$, applied softmax, and then used them to scale up the value vectors $V$. We saw improved performance, which is likely due to the apparent increase in the number of trainable parameters. However, this version of SNN attention had matrix multiplication and softmax bottleneck, and hence was undesirable.
IV-D V-index cache
A naive approach to implement the SNN attention mechanism is to create $n_{inp}\times n_{inp}$ concatenated pairs of the form (14) and then apply the LUT (15) to each of them, resulting in quadratic scaling, $\mathcal{O}(n^{2}_{inp})$.
A better approach, one that achieves linear scaling, is as follows. For simplicity, assume that $V_{z}$ has only one look-up table, so we omit the subscript $i$. We apply the function $H$ (Eq. 1), using $n_{c}$ anchor neurons, to every embedding $z_{\text{pos}}$ to get the look-up row index $j_{\text{pos}}$. We similarly apply $H$, using a different set of $p$ anchor neurons, to every positional encoder $PE_{\text{pos}}$ to get $j^{p}_{\text{pos}}$. This pre-computation is performed once for all $n_{inp}$ positions. We call it V-index caching. Now, to obtain the index for any concatenated vector $z_{\text{pos}_{1}},z_{\text{pos}_{2}},PE_{\text{pos}_{1}-\text{pos}_{2}}]$ needed by the look-up table $V_{z}$, we simply concatenate the three cached indices: $j_{\text{pos}_{1}},j_{\text{pos}_{2}}$, and $j^{p}_{\text{pos}_{1}-\text{pos}_{2}}$. This mechanism is equivalent to the one in Fig. 12b with $z_{i}$, $z_{j}$, and $PE_{i-j}$ being replaced by their respective cached indices.
V-index cache suggests a simple interpretation of how our approach relates to the standard attention mechanism (13): computing an index for $z_{i}$ corresponds to computing the query vector $Q_{i}$. Computing the index for $z_{j}$ corresponds to computing the key vector $K_{j}$. Concatenating them corresponds to the dot-product $Q_{i}K^{\top}_{j}$, and using the concatenated index in the LUT (15) corresponds to multiplying the dot product by the value vector $V_{j}$.
V Experiments
There should be no expectation that SNNs will immediately outperform ANNs because ANNs have benefited from two decades of technological refinement, yielding performance-boosting techniques such as dropout, L2 regularization, or Adam optimization. All these performance-boosting techniques have yet to be invented for SNNs of the form (3).
In contrast, it is reasonable to expect that SNNs will outperform ANNs when compared apples-to-apples—that is, without performance-improving regularization or optimization. To illustrate this point, we use the standard byte-size token prediction task [25] – providing the model with 32-character snippets of English text and requiring it to predict the next character. We performed no hyperparameter search, relying solely on sanity check with a few initial choices of $n$, $n_{t}$, and $n_{c}$, and initialize all synapses to 0. All SNN simulations were performed on author’s 2022 MacBook Air. Simulations of SNN transformer were stopped after its validation perplexity dropped below 2.0 (bits-per-character dropped below 1.00). The code is available at https://github.com/izhikevich/SNN.
V-A Spiking RNN
Let us illustrate a few interesting properties of SNNs using spiking RNN of the form (11), with parameters detailed in Table I.
TABLE I: Parameter values of spiking RNNs (11)
Parameter Value
Context size $n_{inp}$ 32
Dimension of hidden state $h_{t}$ $n$ 64
Memory footprint of embedder table $E$ $256\times n$ 16K
look-up transformation $S_{h}$
Number of look-up tables $n_{t}$ 64
Number of comparisons in each table $n_{c}$ 10
Memory footprint $n_{t}\,2^{n_{c}}n$ 4M
Memory bandwidth per token $2n_{t}n_{c}+n_{t}n$ 5.4K
unembedder look-up transformation $U_{h}$
Number of look-up tables $n_{t}$ 64
Number of comparisons in each table $n_{c}$ 6
Memory footprint $n_{t}\,2^{n_{c}}256$ 1M
Memory bandwidth per token $2n_{t}n_{c}+n_{t}256$ 17K
Size of the model 5M
It might seem counterintuitive in the context of ANNs that the memory bandwidth of the spiking RNN is significantly smaller than its synaptic footprint, since the two are typically the same in ANNs. Let us perform a careful calculation: for each input, $S_{h}$ needs to load $2n_{t}n_{c}$ anchor neurons to determine the indices $j$, and then load $n_{t}$ synaptic vectors $S_{ij}\in\mathbb{R}^{n}$, requiring only 5.4K values. The rest of the 4M synaptic values are entirely passive, becoming relevant only when a new input triggers the generation of different indices $j$. This structure enables efficient scaling: increasing the number of comparisons, $n_{c}$, by 1 doubles the synaptic memory footprint, yet only marginally increases the active computation (FLOPs or memory bandwidth). This unique feature allows SNNs to learn more data without substantially increasing the computational cost, thereby boosting the data-per-FLOP and data-per-bandwidth efficiency ratios.
A notable characteristic of spiking RNNs is their resilience to vanishing or exploding gradients, an advantage that stems from two main architectural features. First, the LUT $S_{h_{t}}$ selects different rows $j$ for distinct hidden states $h_{t}$ during a forward pass, thereby avoiding the repetitive multiplication by the same synaptic matrix that characterizes standard RNNs. Second, although small gradients may lead to small weights and correspondingly small magnitude of hidden state vectors, the selection of the look-up index is based on pairwise comparisons of elements of these vectors. The network’s function is largely unaffected by the absolute magnitude of the components; whether they all have a magnitude of $10^{+10}$ or $10^{-10}$ does not alter the outcome of pairwise comparisons and hence the output to the next layer. The magnitudes, however, affect the learning rate via the uncertainty function $U^{\prime}$ in (8).
The standard metric for measuring byte-size LLM performance is Bits-Per-Character (BPC), which is the cross-entropy loss calculated using the base-2 logarithm rather than the common natural logarithm. As Table II shows, despite its small size and the absence of LSTM gates, regularization, or optimizations, the 5M-parameter spiking RNN would have been competitive with the state-of-the-art LSTMs prior to the invention of Transformers in 2017.
TABLE II: Entropy in bits-per-character (BPC)
Model year size BPC
MI-LSTM [26] 2016 17M 1.44
mLSTM [27] 2016 10M 1.40
Spiking RNN (Eq. 11) 5M 1.39
BN LSTM [28] 2016 16M 1.36
HM-LSTM [29] 2016 35M 1.32
LN HM-LSTM [30] 2016 35M 1.29
SNN transformer (Eq. 12, 15) 806M 0.99
V-B SNN Transformers
We use the same byte-size token prediction task as above to compare the performance of ANN and SNN transformers in Fig. 11. For the ANN transformer, we use exactly the same parameters as in the original paper [24], i.e., $N=6$ layers, $h=8$ heads per layer, $d_{model}=512$, $d_{ff}=2048$, $d_{k}=64$. The ANN simulation, provided by Taylor Kergan from UCSC, was performed without auxiliary training aids such as dropout, Adam optimization, or RMSNorm. Parameters of SNN transformers are provided in Table III.
TABLE III: Parameter values of SNN transformers in Fig. 13
Parameter SNN Attention-only
(green curve) (blue curve)
Context size $n_{inp}$ 32 32
Layers NN 6 6
Embedding dimension $n$ 32 16
Attention $V_{z}$
Heads $h$ 4 1
Look-up tables $n_{t}$ 16 10
Comparisons $n_{c}$ 6 6
Positional dimension pp 4 4
Memory footprint $n_{t}\,2^{2n_{c}+p}n$ 33M 10.5M
Feedforward network $S_{x}$ removed
Look-up tables $n_{t}$ 16 -
Comparisons $n_{c}$ 6 -
Memory footprint $n_{t}\,2^{n_{c}}n$ 33K -
Size of the model 806M 63M
Notice the striking difference in memory allocation between ANN and SNN transformers. In ANNs, the majority of the synaptic memory is consumed by the feedforward network. Conversely, most synaptic memory in SNN transformers is consumed by the attention layer. This allocation is driven by the large size of the pairwise concatenated vectors (in Eq. 14), which necessitates many comparisons and exponentially increases the size of each look-up table to $2^{2n_{c}+p}$. This large memory footprint is the price paid for the greater efficiency of attention in SNN transformers.
We use the parameters of the attention-only SNN transformer to compare resource demands of the ANN and SNN implementations in Table IV. Notice the 10,000-fold difference in memory bandwidth demand: ANNs need to load all $W$ matrices for each forward pass. Batch learning mitigates this issue by spreading the loading cost over many training examples, thereby narrowing the gap with SNNs.
TABLE IV: Resource Demand for Inference per Layer per head in Fig. 11
ANN transformer SNN transformer
Computational cost
$QK^{\top}$ multiplications $2d_{k}n^{2}_{inp}+2d^{2}_{model}n_{inp}$ 0
$QK^{\top}$ summations $2d_{k}n^{2}_{inp}+2d^{2}_{model}n_{inp}$ 0
$V$ and $O$ multiplications $2d_{k}n^{2}_{inp}+4d^{2}_{model}n_{inp}$ 0
$V$ and $O$ summations $2d_{k}n^{2}_{inp}+4d^{2}_{model}n_{inp}$ $n_{t}nn^{2}_{inp}$
FFN multiplications $8d^{2}_{model}n_{inp}$ 0
FFN summations $8d^{2}_{model}n_{inp}$ $n_{t}nn_{inp}$
Comparisons & concatenations 0 $2n_{t}n_{c}n_{inp}$
Total 235,405,312 172,800
Memory footprint
$W^{Q}$, $W^{K}$, $W^{V}$, $W^{O}$ matrices $4d^{2}_{model}$
FFN’s $W^{1}$ and $W^{2}$ matrices $8d^{2}_{model}$
$V_{x}$ look-up transformation $n_{t}n2^{2n_{c}+p}$
$S_{x}$ look-up transformation $n_{t}n2^{n_{c}}$
Total 3,145,728 10,496,000
Attention memory bandwidth demand per new token
$W^{Q}$, $W^{K}$, $W^{V}$, $W^{O}$ matrices $4d^{2}_{model}$
KV-cache $(d_{k}+d_{model})n_{inp}$
Anchor neurons $2n_{t}n_{c}$
V-index cache $3n_{t}n_{inp}$
[…]
Figure 13: Illustration of (a) the 50-fold difference in learning convergence rates and (b) the 10,000-fold difference in required computational resources of ANN and SNN transformers. Training was performed using the Adam learning rate scheduler that was optimized for ANN performance [24]. Notice that the attention-only SNN has 16-dimension embedding space and a single attention head, whereas the ANN transformer has 512-dimensional embedding space and 8 heads.
Figure 13 illustrates another striking difference between ANN and SNN transformers: the learning convergence rate, as measured by performance on a validation dataset during training. This almost 50-fold difference cannot be dismissed as being due to a wrong learning rate parameter, especially since we used the Adam learning rate scheduler with parameters optimized for the ANN transformer, not the SNN one.
Since SNN transformers are built on an architecture fundamentally optimized for efficiency and low memory bandwidth demand, we experimented with reducing the embedding dimension to $n=16$. This configuration still has greater encoding capacity, $16!$, than the linear encoding capacity of $\mathbb{R}^{512}$ – the embedding dimension of ANN transformer, discussed in Sect. VIII-A. To further test the benefits of SNNs, we remove the feedforward layer completely. In this ”attention-only” transformer, plotted as blue curve in the figure, the output from attention module in layer $l$ goes directly to the attention module in layer $l+1$ and the whole architecture is just a stack of $N=6$ blocks in Fig. 12b. Unlike scaled dot-product attention (13) of the ANN transformer, LUTs are inherently nonlinear. Therefore, the SNN attention is all you need to perform both token mixing and channel mixing.
Figure 14: A generalization of the SNN framework presented in this manifesto: Forward pass of state $x$: A locality-sensitive hash function $H$ partitions the input space $X$ into a finite set of buckets $B$, numbered 1, 2, 3, etc. An input $x\in X$ is mapped to the bucket with index $j=H(x)$, which projects to a point $S_{j}\in Y$ in the output space. Consequently, all points within the red bucket project to the same output $S_{j}$, promoting generalization during inference. Backward pass of blue vector $v$: Find the nearest neighbor $\bar{x}$ to $x$ that causes the hash index to flip, resulting in a different (green) bucket $\bar{j}=H(\bar{x})$. This transition moves the output from $S_{j}$ to $S_{\bar{j}}$. We measure the alignment of the resulting vector shift ($S_{\bar{j}}-S_{j}$) with the vector $v$ using a scalar $g$ (e.g., via dot product or projection on $v$). The surrogate (back-propagated) image of $v$ is along the vector $\bar{x}-x$ scaled by $g$ and function $U^{\prime}(|\bar{x}-x|)$ defined in Fig. 8. If $g<0$, the surrogate vector points away from $\bar{x}$.
VI Spiking as Locality-Sensitive Hashing (LSH)
The function (1) that maps latency vectors $x\in\mathbb{R}^{n}$ to integer indices $j$ has a property of LSH (see [15]): nearby vectors $x$ are mapped to the same index with high probability, whereas distant vectors are not. The LSH perspective inspires the generalized framework illustrated in Fig. 14.
  • Forward pass (the hash): The process of partitioning the continuous input space $X$ into a huge but finite number of discrete, non-overlapping hash buckets B using the classic definition of LSH. Each spiking pattern falls into a hash bucket. Small perturbations of input fall into the same bucket, reducing sensitivity to noise, decreasing overfitting, and promoting the generalization in the network.
  • Backward pass (the gradient): Since the hash function is not differentiable, learning relies on surrogate gradients, which are computed by determining the nearest neighbor, $\bar{x}$, of $x$ belonging to a different bucket, $\bar{j}$. Then, the surrogate gradient has the direction of the difference vector, $\bar{x}-x$, scaled by the error alignment $g$.
Our implementation of SNN partitions $X=\mathbb{R}^{n}$ into $2^{n_{t}n_{c}}$ discrete buckets. The output of each bucket, $S_{j}$ is defined by (3). Learning is just a nudge of $S_{j}$ along vector $v$. Error backpropagation is achieved by identifying the minimal input perturbation $\bar{x}$ that flips the relative spike order of a single anchor pair, $u_{i}$, in each table $i$. The resulting output shift ($S_{i\bar{j}}-S_{ij}$) is used in (7) to measure the alignment, $g_{i}$, with the error vector $v$. The surrogate gradient (8) is the average of all such perturbations $(\bar{x}-x)$ scaled by their alignments $g_{i}$ and smoothened by the function $U(|\bar{x}-x|)$ defined in Fig. 8. This method efficiently samples the error signal along the most relevant directions determined by the look-up table boundaries.
Observe that matrix multiplication is not required for the forward pass; it is, however, required for the backward pass to calculate the alignment $g_{i}$ for each look-up table. A method of finding $g_{i}$ using a LUT of the form (3) would eliminate MatMul operations and the need for GPUs during training.
In this general formulation, we have freedom of choosing the hash function $H(x)$. Useful functions are those requiring few computations, producing large partitions, allowing simple look-up table representations, and readily rendering $\bar{x}$ for any $x$. Our choice (1) satisfies all these requirements. Other choices are discussed in Sect. VIII-C, and it seems that any choice inspired by spike-timing patterns leads to a useful $H(x)$.
VII Conclusion
While both ANNs and SNNs rely on vectors, the interpretation of those vectors is what truly dictates how we manipulate them. For ANNs, the pairwise relationships between vectors—such as orthogonality and correlation—are paramount. For SNNs, however, the relationships among elements within each vector are paramount, as these relationships correspond to the different timing of spikes. This difference in interpretation leads to fundamentally different computational approaches. However, SNNs retain the most important operational aspect of ANNs: learning by gradient descent. This retention allows us to successfully reproduce deep networks, RNNs, and transformers, opening the possibility to integrate other architectural innovations developed for ANNs into SNNs.
Acknowledgement: Oleg Sinavski and Mikhail Burtsev read the first draft of this manifesto and made me rewrite it completely. Anatoli Starostin, Frank C. Hoppensteadt, Alfredo Maria Solano, Takeshi Kojima, Kan Hatakeyama, Assaf Touboul, Ilya Rybak, Bill MacCartney, Thomas Stahura, and Todd Hylton have made a number of useful suggestions. Special thanks to Taylor Kergan from UCSC, who provided simulation results of the ANN transformer.